Suppose ABC corp. is evaluating a project with the following incremental free-cash-flows (FCF) paid at the end of the next four years. Which annual discount rate makes the NPV equal to zero? Enter your answer as a percent do not include the %. Round your final answer to two decimals.

Timeline

0

1

2

3

4

FCF

-500

200

200

200

200
Answer #1

Internal rate of return is the the annual discount rate that makes NPV equal to 0

NPV = Present value of cash inflows - present value of cash outlows

NPV = -500 + 200 / (1 + R)1 + 200 / (1 + R)2 + 200 / (1 + R)3 + 200 / (1 + R)4

Using trial and error method, i.e., after trying various values for R, lets try R as 21.86%

NPV = -500 + 200 / (1 + 0.2186)1 + 200 / (1 + 0.2186)2 + 200 / (1 + 0.2186)3 + 200 / (1 + 0.2186)4

NPV = -500 + 164.1228 + 134.6814 + 110.5214 + 90.6954

NPV = 0

Therefore, Annual discount rate is 21.86%